OpenStudy (gabbyalicorn):

Which answer shows this equation written in standard form? 4 – 3(x – y) = –8x + 2 A. –11x + 3y = –2 B. 5x – 3y = 2 C. 5x + 3y = –2 D. 5x + y = 2

OpenStudy (gabbyalicorn):

I don't get this at all D:

OpenStudy (er.mohd.amir):

5x+3y=-2

OpenStudy (dancergirl45):

Hi there! openstudy is a place to answer questions while helping the asker learn how to solve the problem, not do their work for them. Hope you understand! Thankies!

OpenStudy (gabbyalicorn):

OMG u copied my thing from PS

OpenStudy (dancergirl45):

yeah see how rude it is

OpenStudy (gabbyalicorn):

>:I its not rude

OpenStudy (ihelpyoulearn):

I agree with @Er.Mohd.AMIR

OpenStudy (jango_in_dtown):

you have to write in ax+by=c form 4 – 3(x – y) = –8x + 2 implies -3x+3y=-8x+2-4 implies -3x+8x+3y=-2 implies 5x+3y=-2

OpenStudy (gabbyalicorn):

srsly i didnt do anything wrong @dancergirl45 so why would you be saying that to me

OpenStudy (dancergirl45):

ok then hope no offense was taken :p

OpenStudy (gabbyalicorn):

Thank you @Er.Mohd.AMIR and @IHelpYouLearn

OpenStudy (gabbyalicorn):

Also, @dancergirl45 since you blocked me I was going to say i only said that because the person wanted to do their whole assignment for them.